OpenStudy (anonymous):

Finish these three solutions to this equation. Y= x-1/2

OpenStudy (anonymous):

( ,1/2,) ( ,3/2) ( ,5/2) i forgot to give this sorry

OpenStudy (amorfide):

you are given the y values substitute y=1/2 make x the subject substitute y=3/2 do the same y=5/2 do the same

OpenStudy (anonymous):

how do i do it i dont know how to do it.

OpenStudy (anonymous):

@Sir_Rico_of_Eureka can u help nobody explain how to do this if someone would explain how to get one i can get the others can u plz help me

OpenStudy (anonymous):

If you think about each set of numbers as (x,y) It will be easier to visualize it. \[y=x-\frac{ 1 }{ 2}\] Set y equal to 1/2. 3/2. and 5/2. You'll get 1/2 = x- 1/2, 3/2=x-1/2 and 5/2=x-1/2

OpenStudy (anonymous):

Can you solve the rest?

OpenStudy (anonymous):

i get how to do all that it's just i dont know how to get the answer

OpenStudy (anonymous):

You want to get x on its own. Subtract \[\frac{ 1 }{ 2 } \] from boths sides. For all the equations. \[\frac{ 1 }{ 2 } = x-\frac{ 1 }{ 2 }\] \[\frac{ 3 }{ 2 } = x -\frac{ 1 }{ 2 }\] \[\frac{ 5 }{ 2 } = x - \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

so the 1/2n one would be zero right?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

OHHHHH NOOOO!

OpenStudy (anonymous):

I meant ADD! I'm soooo sorry.

OpenStudy (anonymous):

what? lol

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

so 1/2 would = 2/2=1

OpenStudy (anonymous):

:]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

3/2=6/2=3

OpenStudy (anonymous):

3/2 + 1/2 = 4/2 = 2

OpenStudy (anonymous):

sorry add wrong lol

OpenStudy (anonymous):

5/2=6/2=3

OpenStudy (anonymous):

thank you for ur help i see how to do it now